\begin{thebibliography}{99}
\bibitem{Br1} A. Brannan and J. G. Clunie, Aspects of contemporary complex
analysis Proceedings of the NATO Advanced Study Institute held at the
University of Durham, Durham, July 120, 1979, Academic Press New York,
London, 1980.

\bibitem{Bran} D.A. Brannan and T.S. Taha, On some classes of bi-univalent
functions, in: S.M. Mazhar, A. Hamoui, N.S. Faour (Eds.), Math. Anal. and
Appl., Kuwait; February 18--21, 1985, in: KFAS Proceedings Series, vol. 3,
Pergamon Press, Elsevier Science Limited, Oxford, 1988, pp. 53--60. see also
Studia Univ. Babeş-Bolyai Math. 31 (2) (1986), 70--77.

\bibitem{Bul} S. Bulut, Faber polynomial coefficient estimates for a
comprehensive subclass of analytic bi-univalent functions, C. R. Acad. Sci.
Paris, Ser. I, 352 (6) (2014), pp. 479--484.

\bibitem{Cantor} D. G. Cantor, Power series with integral coefficients,
Bull. Amer. Math. Soc. 69 (1963), 362--366.

\bibitem{CaOrNi} M. Ça\u{g}lar, H. Orhan and N. Ya\u{g}mur, Coefficient
bounds for new subclasses of bi-univalent functions, Filomat 27:7 (2013),
1165-1171.

\bibitem{Deniz} E. Deniz, Certain subclasses of bi-univalent functions
satisfying subordinate conditions, J. Class. Anal. 2(1) (2013), 49--60.

\bibitem{Duren 1} P. L. Duren, Univalent functions,Grundlehren der
Mathematischen Wissenschaften, 259, Springer, New York, 1983.

\bibitem{Fekete} M. Fekete and G. Szegö, Eine Bemerkung uber ungerade
schlichte Funktionen, J. London Math. Soc. 8 (1933), 85-89.

\bibitem{Fra} B.A. Frasin, M.K. Aouf, New subclasses of bi-univalent
functions, Appl. Math. Lett. 24 (2011) 1569-1573.

\bibitem{Grenan} U. Grenander and G. Szegö, Toeplitz forms and their
applications, California Monographs in Mathematical Sciences Univ.
California Press, Berkeley, 1958.

\bibitem{Jahan} S. G. Hamidi, J. M. Jahangiri, Faber polynomial coefficient
estimates for analytic bi-close-to-convex functions, C. R. Acad. Sci. Paris,
Ser. I 352(2014) 17--20.

\bibitem{Ja} A, Janteng, S. A. Halim and M. Darus, Hankel Determinant for
starlike and convex functions. Int. J. Math. Anal. 1(13) (2007), 619-625.

\bibitem{Ked} A. W. Kedzierawski, Some remarks on bi-univalent functions,
Ann. Univ.Mariae Curie-Skłodowska Sect. A 39 (1985), 77--81 (1988).

\bibitem{ke} F. R. Keogh and E. P. Merkes, A coefficient inequality for
certain classes of analytic functions, Proc. Amer. Math. Soc. 20 (1969),
8--12.

\bibitem{Ku} S. S. Kumar, V. Kumar and V. Ravichandran, Estimates for the
initial coefficients of bi-univalent functions, Tamsui Oxford J. Inform.
Math. Sci. 29(4) (2013), 487-504.

\bibitem{Lee} S. K. Lee, V. Ravichandran and S. Supramaniam, Bounds for the
second Hankel determinant of certain univalent functions, J. Ineq. Appl.
2013, 2013:281.

\bibitem{Lew} M. Lewin, On a coefficient problem for bi-univalent functions,
Proc. Amer. Math. Soc. 18 (1967), 63--68.

\bibitem{Ne} E. Netanyahu, The minimal distance of the image boundary from
the origin and the second coefficient of a univalent function in $|z|<1$,
Arch. Rational Mech. Anal. 32 (1969), 100--112.

\bibitem{No-Tho} J.W. Noonan and D.K. Thomas, On the second Hankel
determinant of areally mean p-valent functions,Trans. Amer. Math. Soc.,
223(2) (1976), 337-346.

\bibitem{Or} H. Orhan, E. Deniz and D. Raducanu, The Fekete Szegö problem
for subclasses of analytic functions defined by a differential operator
related to conic domains, Comput. Math. Appl. 59 (2010), 283-295.

\bibitem{Or2} H. Orhan, N. Magesh, and V. K. Balaji, Initial coefficient
bounds for a general class of bi-univalent functions, Filomat, (2014) in
press.

\bibitem{Pom} Ch. Pommerenke, Univalent Functions, Vandenhoeck and Rupercht,
Göttingen, 1975.

\bibitem{Sri} H. M. Srivastava, A.K. Mishra, P. Gochhayat, Certain
subclasses of analytic and bi-univalent functions, Appl. Math. Lett. 23
(2010) 1188--1192.

\bibitem{Sri2} H. M. Srivastava, S. Bulut, M. Ça\u{g}lar, N. Ya\u{g}mur,
Coefficient estimates for a general subclass of analytic and bi-univalent
functions, Filomat 27:5 (2013), 831-842.

\bibitem{Tan} D. L. Tan, Coefficient estimates for bi-univalent functions,
Chinese Ann. Math. Ser. A 5 (5) (1984), 559--568.

\bibitem{Qing1} Q.-H. Xu, Y.-C. Gui and H. M. Srivastava, Coefficient
estimates for a certain subclass of analytic and bi-univalent functions,
Appl. Math. Lett. 25 (2012) 990--994.

\bibitem{Qing2} Q.-H. Xu, H.-G. Xiao and H. M. Srivastava, A certain general
subclass of analytic and bi-univalent functions and associated coefficient
estimate problems, Appl. Math. Comput. 218 (2012), 11461-11465.

\bibitem{Za} P. Zaprawa, On the Fekete-Szegö problem for classes of
bi-univalent functions, Bull. Belg. Math. Soc. Simon Stevin 21 (2014),
169--178.

\bibitem{Za1} P. Zaprawa, Estimates of initial coefficients for bi-univalent
functions, Abstr. Appl. Anal., 2014, Article ID 357480, 6 pages.
\end{thebibliography}

\documentclass[a4paper,10pt,leqno]{amsart}
\usepackage[cp1250]{inputenc}
\usepackage{ifthen}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{latexsym}

\numberwithin{equation}{section}
\newtheorem{theorem}{Theorem}[section]
\theoremstyle{plain}

\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{lemma}{Lemma}[section]

\numberwithin{equation}{section}
\input{tcilatex}

\begin{document}
\title[Second Hankel determinant for bi-starlike and bi-convex functions of
order $\beta $]{Second Hankel determinant for bi-starlike and bi-convex
functions of order $\beta $}
\author{Erhan Deniz}
\author{Murat Ça\u{g}lar}
\address{Department of Mathematics, Faculty of Science and Letters, Kafkas
University, Kars, Turkey.}
\email{edeniz36@gmail.com (Erhan Deniz), mcaglar25@gmail.com (Murat Ça\u{g}%
lar)}
\author{Halit Orhan}
\address{Department of Mathematics, Faculty of Science, Ataturk University,
Erzurum, 25240, Turkey.}
\email{orhanhalit607@gmail.com (Halit Orhan)}
\subjclass[2000]{ Primary 30C45, 30C50; Secondary 30C80.}
\keywords{Bi-univalent functions, bi-starlike functions of order $\beta $,
bi-convex functions of order $\beta $, second Hankel determinant. \\
\textit{Corresponding author}. edeniz36@gmail.com (Erhan Deniz)}

\begin{abstract}
In the present investigation the authors obtain upper bounds for the second
Hankel determinant $H_{2}(2)$ of the classes bi-starlike and bi-convex
functions of order $\beta $, represented by $S_{\sigma }^{\ast }(\beta )\;$%
and $K_{\sigma }(\beta )$, respectively. In particular, the estimates for
the second Hankel determinant $H_{2}(2)$ of bi-starlike and bi-convex
functions which are important subclasses of bi-univalent functions are
pointed out.
\end{abstract}

\maketitle

\section{Introduction and definitions}

Let $\mathcal{A}$ denote the family of functions $f$\ analytic in the open
unit disk $\mathcal{U}=\left\{ {z\in 
\mathbb{C}
:\left\vert z\right\vert <1}\right\} $ of the form 
\begin{equation}
f(z)=z+\sum\limits_{n=2}^{\infty }{a_{n}z^{n}.}  \label{eq1}
\end{equation}%
Let $\mathcal{S}$ denote the class of all functions in $\mathcal{A}$ which
are univalent in $\mathcal{U}$. The Koebe one-quarter theorem (see \cite%
{Duren 1}) ensures that the image of $\mathcal{U}$ under every $f\in 
\mathcal{S}$ contain a disk of radius $1\diagup 4.$ So, every $f\in \mathcal{%
S}$ has an inverse function $f^{-1}$ satisfying $f^{-1}(f(z))=z$ $\left(
z\in \mathcal{U}\right) $ and%
\begin{equation*}
f(f^{-1}(w))=w\text{ \ \ }\left( \left\vert w\right\vert <r_{0}(f);\text{ }%
r_{0}(f)\geq 1\diagup 4\right)
\end{equation*}%
where $%
f^{-1}(w)=w-a_{2}w^{2}+(2a_{2}^{2}-a_{3})w^{3}-(5a_{2}^{3}-5a_{2}a_{3}+a_{4})w^{4}+.... 
$

A function $f\in \mathcal{A}$ is said to be bi-univalent in $\mathcal{U}$ if
both $f(z)$ and $f^{-1}(z)$ are univalent in $\mathcal{U}.$ Let $\sigma $
denote the class of bi-univalent functions in $\mathcal{U}$ given by (\ref%
{eq1}).

Two of the most famous subclasses of univalent functions are the class $%
\mathcal{S}^{\ast }(\beta )$ of starlike functions of order $\beta $ and the
class $\mathcal{K}(\beta )\;$of convex functions of order $\beta $. By
definition, we have%
\begin{equation*}
\mathcal{S}^{\ast }(\beta )=\left\{ f\in \mathcal{S}:\Re \left( \frac{%
zf^{\prime }(z)}{f(z)}\right) >\beta ;\;z\in \mathcal{U};\;0\leq \beta
<1\right\}
\end{equation*}%
and%
\begin{equation*}
\mathcal{K}(\beta )=\left\{ f\in \mathcal{S}:\Re \left( 1+\frac{zf^{\prime
\prime }(z)}{f^{\prime }(z)}\right) >\beta ;\;z\in \mathcal{U};\;0\leq \beta
<1\right\} .
\end{equation*}%
The classes consisting of starlike and convex functions are usually denoted
by $\mathcal{S}^{\ast }=\mathcal{S}^{\ast }(0)$ and $\mathcal{K}=\mathcal{K}%
(0)$, respectively.

For $0\leq \beta <1,$ a function $f\in \sigma $ is in the class $\mathcal{S}%
_{\sigma }^{\ast }(\beta )$ of \textit{bi-starlike functions of order} $%
\beta ,$ or $\mathcal{K}_{\sigma }(\beta )$ of \textit{bi-convex functions
of order} $\beta $ if both $f$ and its inverse map $f^{-1}$ are,
respectively, starlike or convex of order $\beta .$ These classes were
introduced by Brannan and Taha \cite{Bran} in 1985. Especially the classes $%
\mathcal{S}_{\sigma }^{\ast }(0)=\mathcal{S}_{\sigma }^{\ast }$ and $%
\mathcal{K}_{\sigma }(0)=\mathcal{K}_{\sigma }$ are \textit{bi-starlike} and 
\textit{bi-convex functions}, respectively. In 1967, Lewin \cite{Lew} showed
that for every functions $f\in \sigma $ of the form (\ref{eq1}), the second
coefficient of $f$ satisfy the inequality $|a_{2}|<1.51$. In 1967, Brannan
and Clunie \cite{Br1} conjectured that $|a_{2}|\leq \sqrt{2}$ for $f\in
\sigma .$ Later, Netanyahu \cite{Ne} proved that $\max_{f\in \sigma
}|a_{2}|=4/3$. \ In 1985, Kedzierawski \cite{Ked} proved Brannan and
Clunie's conjecture for $f\in \mathcal{S}_{\sigma }^{\ast }$. In 1985, Tan 
\cite{Tan} obtained the bound for $a_{2}$ namely $|a_{2}|<1.485$ which is
the best known estimate for functions in the class $\sigma .$ Brannan and
Taha \cite{Bran} obtained estimates on the initial coefficients $\left\vert
a_{2}\right\vert $ and $\left\vert a_{3}\right\vert $ for functions in the
classes $\mathcal{S}_{\sigma }^{\ast }(\beta )$ and $\mathcal{K}_{\sigma
}(\beta )$. Recently, Deniz \cite{Deniz} and Kumar et al. \cite{Ku} both
extended and improved the results of Brannan and Taha \cite{Bran} by
generalizing their classes using subordination. The problem of estimating
coefficients $\left\vert a_{n}\right\vert $, $n\geq 2$ is still open.
However, a lot of results for $\left\vert a_{2}\right\vert $, $\left\vert
a_{3}\right\vert $ and $\left\vert a_{4}\right\vert $ were proved for some
subclasses of $\sigma $ (see \cite{Bul}, \cite{CaOrNi}, \cite{Fra}, \cite%
{Jahan}, \cite{Or2}, \cite{Sri}, \cite{Sri2}, \cite{Qing1}, \cite{Qing2}).
Unfortunatelly, none of them are not sharp.

One of the important tools in the theory of univalent functions is Hankel
Determinants which are utility, for example, in showing that a function of
bounded characteristic in $\mathcal{U}$, i.e., a function which is a ratio
of two bounded analytic functions, with its Laurent series around the origin
having integral coefficients, is rational \cite{Cantor}. The Hankel
determinants \cite{No-Tho}$\;H_{q}(n)\;(n=1,2,...,\;q=1,2,...)$ of the
function $f$\ are defined by%
\begin{equation*}
H_{q}(n)=\left[ 
\begin{array}{cccc}
a_{n} & a_{n+1} & ... & a_{n+q-1} \\ 
a_{n+1} & a_{n+2} & ... & a_{n+q} \\ 
\vdots & \vdots &  & \vdots \\ 
a_{n+q-1} & a_{n+q} & ... & a_{n+2q-2}%
\end{array}%
\right] \text{ \ \ }(a_{1}=1).
\end{equation*}

This determinant was discussed by several authors with $q=2$. For example,
we can know that the functional $H_{2}(1)=a_{3}-a_{2}^{2}\;$is known as the
Fekete-Szegö functional and they consider the further generalized functional 
$a_{3}-\mu a_{2}^{2}\;$where $\mu $ is some real number (see, \cite{Fekete}%
). In 1969, Keogh and Merkes \cite{ke} proved the Fekete-Szegö problem for
the classes $\mathcal{S}^{\ast }$ and $\mathcal{K}$. Someone can see the
Fekete-Szegö problem for the classes $\mathcal{S}^{\ast }(\beta )$ and $%
\mathcal{K}(\beta )$ at special cases in the paper of Orhan \textit{et.al. }%
\cite{Or}. On the other hand, very recently Zaprawa \cite{Za}, \cite{Za1}
have studied on Fekete-Szegö problem for some classes of bi-univalent
functions. In special cases, he gave Fekete-Szegö problem for the classes $%
\mathcal{S}_{\sigma }^{\ast }(\beta )$ and $\mathcal{K}_{\sigma }(\beta )$.
In 2014, Zaprawa \cite{Za} proved the following resuts for $\mu \in 
\mathbb{R}
,$%
\begin{equation*}
f\in \mathcal{S}_{\sigma }^{\ast }(\beta )\Rightarrow \left\vert a_{3}-\mu
a_{2}^{2}\right\vert \leq \left\{ 
\begin{array}{cc}
1-\beta ; & \frac{1}{2}\leq \mu \leq \frac{3}{2} \\ 
2(1-\beta )\left\vert \mu -1\right\vert ; & \mu \geq \frac{3}{2}\text{ and }%
\mu \leq \frac{1}{2}%
\end{array}%
\right.
\end{equation*}%
and%
\begin{equation*}
f\in \mathcal{K}_{\sigma }(\beta )\Rightarrow \left\vert a_{3}-\mu
a_{2}^{2}\right\vert \leq \left\{ 
\begin{array}{cc}
\frac{1-\beta }{3}; & \frac{2}{3}\leq \mu \leq \frac{4}{3} \\ 
(1-\beta )\left\vert \mu -1\right\vert ; & \mu \geq \frac{4}{3}\text{ and }%
\mu \leq \frac{2}{3}%
\end{array}%
\right. .
\end{equation*}

The second Hankel determinant $H_{2}(2)$ is given by $%
H_{2}(2)=a_{2}a_{4}-a_{3}^{2}.$ The bounds for the second Hankel determinant 
$H_{2}(2)$ obtained for the classes $\mathcal{S}^{\ast }$ and $\mathcal{K}$
in \cite{Ja}. Recently, Lee et al. \cite{Lee} established the sharp bound to 
$\left\vert H_{2}(2)\right\vert $ by generalizing their classes using
subordination. In their paper, one can find the sharp bound to $\left\vert
H_{2}(2)\right\vert $ for the functions in the classes $\mathcal{S}^{\ast
}(\beta )$ and $\mathcal{K}(\beta )$.

In this paper, we seek upper bound for the functional $%
H_{2}(2)=a_{2}a_{4}-a_{3}^{2}\;$for functions $f\;$belonging to the classes $%
\mathcal{S}_{\sigma }^{\ast }(\beta )\;$and $\mathcal{K}_{\sigma }(\beta )$.

Let $\mathcal{P}$ be the class of functions with positive real part
consisting of all analytic functions $\mathcal{P}:\mathcal{U\rightarrow 
\mathbb{C}
}$ satisfying $p(0)=1$ and $\Re p(z)>0$.

To establish our main results, we shall require the following lemmas.

\begin{lemma}
\label{l1}\cite{Pom} If the function $p\in \mathcal{P}$ is given by the
series 
\begin{equation}
p(z)=1+c_{1}z+c_{2}z^{2}+...  \label{eq1*}
\end{equation}%
then the sharp estimate $\left\vert c_{k}\right\vert \leq 2$ $(k=1,2,...)$
holds.
\end{lemma}

\begin{lemma}
\label{l2}\cite{Grenan} If the function $p\in \mathcal{P}$ is given by the
series (\ref{eq1*}), then%
\begin{eqnarray}
2c_{2} &=&c_{1}^{2}+x(4-c_{1}^{2})  \label{eq2} \\
4c_{3}
&=&c_{1}^{3}+2(4-c_{1}^{2})c_{1}x-c_{1}(4-c_{1}^{2})x^{2}+2(4-c_{1}^{2})%
\left( 1-\left\vert x\right\vert ^{2}\right) z,  \label{eq3}
\end{eqnarray}%
for some $x,$ $z$ with $\left\vert x\right\vert \leq 1$ and $\left\vert
z\right\vert \leq 1.$
\end{lemma}

\section{Main results}

Our first main result for the class $\mathcal{S}_{\sigma }^{\ast }(\beta )$
as follows:

\begin{theorem}
\label{t1} Let $f(z)$ given by (\ref{eq1}) be in the class $\mathcal{S}%
_{\sigma }^{\ast }(\beta ),$ $0\leq \beta <1.$ Then%
\begin{equation}
\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert \leq \left\{ 
\begin{array}{cc}
\frac{4}{3}\left( 1-\beta \right) ^{2}\left( 4\beta ^{2}-8\beta +5\right) ,
& \beta \in \left[ 0,\frac{29-\sqrt{137}}{32}\right] \\ 
\left( 1-\beta \right) ^{2}\left( \frac{13\beta ^{2}-14\beta -7}{16\beta
^{2}-26\beta +5}\right) , & \beta \in \left( \frac{29-\sqrt{137}}{32}%
,1\right) .%
\end{array}%
\right.  \label{eq5}
\end{equation}
\end{theorem}

\begin{proof}
Let $f\in \mathcal{S}_{\sigma }^{\ast }(\beta )\;$and $g=f^{-1}.$\ Then%
\begin{equation}
\frac{zf^{\prime }(z)}{f(z)}=\beta +(1-\beta )p(z)\;\text{and }\frac{%
wg^{\prime }(w)}{g(w)}=\beta +(1-\beta )q(w)  \label{eq6}
\end{equation}%
where $p(z)=1+c_{1}z+c_{2}z^{2}+...$ and $q(w)=1+d_{1}w+d_{2}w^{2}+...$ in $%
\mathcal{P}$.

Comparing coefficients in (\ref{eq6}), we have%
\begin{eqnarray}
a_{2} &=&(1-\beta )c_{1},  \label{eq8} \\
2a_{3}-a_{2}^{2} &=&(1-\beta )c_{2},  \label{eq9} \\
3a_{4}-3a_{3}a_{2}+a_{2}^{3} &=&(1-\beta )c_{3}  \label{eq10}
\end{eqnarray}%
and%
\begin{eqnarray}
-a_{2} &=&(1-\beta )d_{1},  \label{eq8*} \\
3a_{2}^{2}-2a_{3} &=&(1-\beta )d_{2},  \label{eq9*} \\
-10a_{2}^{3}+12a_{3}a_{2}-3a_{4} &=&(1-\beta )d_{3}.  \label{eq10*}
\end{eqnarray}%
From (\ref{eq8}) and (\ref{eq8*}), we arrive at 
\begin{equation}
c_{1}=-d_{1}  \label{eq11}
\end{equation}%
and%
\begin{equation}
a_{2}=(1-\beta )c_{1}.  \label{eq12}
\end{equation}

Now, from (\ref{eq9}), (\ref{eq9*}) and (\ref{eq12}), we get that%
\begin{equation}
a_{3}=\left( 1-\beta \right) ^{2}c_{1}^{2}+\frac{\left( 1-\beta \right) }{4}%
\left( c_{2}-d_{2}\right) .  \label{eq13}
\end{equation}

Also, from (\ref{eq10}) and (\ref{eq10*}), we find that%
\begin{equation}
a_{4}=\frac{2}{3}\left( 1-\beta \right) ^{3}c_{1}^{3}+\frac{5}{8}\left(
1-\beta \right) ^{2}c_{1}\left( c_{2}-d_{2}\right) +\frac{1}{6}(1-\beta
)\left( c_{3}-d_{3}\right) .  \label{eq14}
\end{equation}%
Thus, we can easily establish that%
\begin{eqnarray}
&&\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert =\left\vert -\frac{1}{3}\left(
1-\beta \right) ^{4}c_{1}^{4}+\frac{1}{8}\left( 1-\beta \right)
^{3}c_{1}^{2}\left( c_{2}-d_{2}\right) \right.  \notag \\
&&\text{ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }\left. +\frac{1}{6}\left( 1-\beta
\right) ^{2}c_{1}\left( c_{3}-d_{3}\right) -\frac{1}{16}\left( 1-\beta
\right) ^{2}\left( c_{2}-d_{2}\right) ^{2}\right\vert .  \label{eq15}
\end{eqnarray}%
According to Lemma \ref{l2} and (\ref{eq11}), we write%
\begin{equation}
\left. 
\begin{array}{c}
2c_{2}=c_{1}^{2}+x(4-c_{1}^{2}) \\ 
2d_{2}=d_{1}^{2}+x(4-d_{1}^{2})%
\end{array}%
\right\} \Longrightarrow c_{2}-d_{2}=\frac{4-c_{1}^{2}}{2}(x-y)  \label{eq16}
\end{equation}%
and%
\begin{eqnarray*}
4c_{3}
&=&c_{1}^{3}+2(4-c_{1}^{2})c_{1}x-c_{1}(4-c_{1}^{2})x^{2}+2(4-c_{1}^{2})%
\left( 1-\left\vert x\right\vert ^{2}\right) z, \\
4d_{3}
&=&d_{1}^{3}+2(4-d_{1}^{2})d_{1}y-d_{1}(4-d_{1}^{2})y^{2}+2(4-d_{1}^{2})%
\left( 1-\left\vert y\right\vert ^{2}\right) w,
\end{eqnarray*}%
\begin{equation}
c_{3}-d_{3}=\frac{c_{1}^{3}}{2}+\frac{c_{1}\left( 4-c_{1}^{2}\right) }{2}%
(x+y)-\frac{c_{1}\left( 4-c_{1}^{2}\right) }{2}(x^{2}+y^{2})+\frac{\left(
4-c_{1}^{2}\right) }{2}\left( \left( 1-\left\vert x\right\vert ^{2}\right)
z-\left( 1-\left\vert y\right\vert ^{2}\right) w\right) .  \label{eq17}
\end{equation}%
for some $x,y,$ $z,w$ with $\left\vert x\right\vert \leq 1,\left\vert
y\right\vert \leq 1,\left\vert z\right\vert \leq 1$ and $\left\vert
w\right\vert \leq 1.$Using (\ref{eq16}) and (\ref{eq17}) in (\ref{eq15}),
and applying the triangle inequality we have%
\begin{eqnarray*}
&&\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert =\left\vert -\frac{1}{3}\left(
1-\beta \right) ^{4}c_{1}^{4}+\frac{1}{16}\left( 1-\beta \right)
^{3}c_{1}^{2}(4-c_{1}^{2})(x-y)\right. \\
&&+\frac{1}{6}\left( 1-\beta \right) ^{2}c_{1}\left[ \frac{c_{1}^{3}}{2}+%
\frac{(4-c_{1}^{2})c_{1}}{2}(x+y)-\frac{(4-c_{1}^{2})c_{1}}{4}(x^{2}+y^{2})+%
\frac{(4-c_{1}^{2})}{2}\left( (1-\left\vert x\right\vert
^{2})z-(1-\left\vert y\right\vert ^{2})w\right) \right] \\
&&\left. -\frac{1}{64}\left( 1-\beta \right)
^{2}(4-c_{1}^{2})^{2}(x-y)^{2}\right\vert \\
&\leq &\frac{1}{3}\left( 1-\beta \right) ^{4}c_{1}^{4}+\frac{1}{12}\left(
1-\beta \right) ^{2}c_{1}^{4}+\frac{1}{6}\left( 1-\beta \right)
^{2}c_{1}(4-c_{1}^{2}) \\
&&+\left[ \frac{1}{16}\left( 1-\beta \right) ^{3}c_{1}^{2}(4-c_{1}^{2})+%
\frac{1}{12}\left( 1-\beta \right) ^{2}c_{1}^{2}(4-c_{1}^{2})\right]
(\left\vert x\right\vert +\left\vert y\right\vert ) \\
&&+\left[ \frac{1}{24}\left( 1-\beta \right) ^{2}c_{1}^{2}(4-c_{1}^{2})-%
\frac{1}{12}\left( 1-\beta \right) ^{2}c_{1}(4-c_{1}^{2})\right] (\left\vert
x\right\vert ^{2}+\left\vert y\right\vert ^{2})+\frac{1}{64}\left( 1-\beta
\right) ^{2}(4-c_{1}^{2})^{2}(\left\vert x\right\vert +\left\vert
y\right\vert )^{2}.
\end{eqnarray*}

Since $p\in $ $\mathcal{P},$ so $\left\vert c_{1}\right\vert \leq 2.$
Letting $c_{1}=c,$ we may assume without restriction that $c\in \lbrack
0,2]. $ Thus, for $\lambda =\left\vert x\right\vert \leq 1$ and $\mu
=\left\vert y\right\vert \leq 1$ we obtain%
\begin{equation}
\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert \leq T_{1}+T_{2}(\lambda +\mu
)+T_{3}(\lambda ^{2}+\mu ^{2})+T_{4}(\lambda +\mu )^{2}=F(\lambda ,\mu ) 
\notag
\end{equation}%
where%
\begin{eqnarray*}
T_{1} &=&T_{1}(c)=\frac{\left( 1-\beta \right) ^{2}}{12}\left[ \left(
1+4\left( 1-\beta \right) ^{2}\right) c^{4}-2c^{3}+8c\right] \geq 0,\text{ }
\\
T_{2} &=&T_{2}(c)=\frac{1}{48}\left( 1-\beta \right)
^{2}c^{2}(4-c^{2})(7-3\beta )\geq 0, \\
T_{3} &=&T_{3}(c)=\frac{1}{24}\left( 1-\beta \right) ^{2}c(4-c^{2})(c-2)\leq
0,\text{ } \\
T_{4} &=&T_{4}(c)=\frac{1}{64}\left( 1-\beta \right)
^{2}(4-c_{1}^{2})^{2}\geq 0.
\end{eqnarray*}

Now we need to maximize $F(\lambda ,\mu )$ in the closed square $\mathbb{S}%
=\left\{ (\lambda ,\mu ):0\leq \lambda \leq 1,0\leq \mu \leq 1\right\} .$
Since $T_{3}<0$ and $T_{3}+2T_{4}>0$ for $c\in \lbrack 0,2)$, we conclude
that%
\begin{equation*}
F_{\lambda \lambda }\cdot F_{\mu \mu }-\left( F_{\lambda \mu }\right) ^{2}<0.
\end{equation*}

Thus the function $F$ cannot have a local maximum in the interior of the
square $\mathbb{S}$. Now, we investigate the maximum of $F$ on the boundary
of the square $\mathbb{S}$.

For $\lambda =0$ and $0\leq \mu \leq 1$ $\left( \text{similarly }\mu =0\text{
and }0\leq \lambda \leq 1\right) ,$ we obtain%
\begin{equation*}
F(0,\mu )=G(\mu )=\left( T_{3}+T_{4}\right) \mu ^{2}+T_{2}\mu +T_{1}.
\end{equation*}

i. \textit{The case }$T_{3}+T_{4}\geq 0:$ In this case for $0<\mu <1$ and
any fixed $c$ with $0\leq c<2,$ it is clear that $G^{\prime }(\mu )=2\left(
T_{3}+T_{4}\right) \mu +T_{2}>0,$ that is, $G(\mu )$ is an increasing
function. Hence, for fixed $c\in \lbrack 0,2),$ the maximum of $G(\mu )$
occurs at $\mu =1,$ and%
\begin{equation*}
\max G(\mu )=G(1)=T_{1}+T_{2}+T_{3}+T_{4}.
\end{equation*}

ii. \textit{The case }$T_{3}+T_{4}<0:$ Since $T_{2}+2\left(
T_{3}+T_{4}\right) \geq 0$ for $0<\mu <1$ and any fixed $c$ with $0\leq c<2,$
it is clear that $T_{2}+2\left( T_{3}+T_{4}\right) <2\left(
T_{3}+T_{4}\right) \mu +T_{2}<T_{2}$ and so $G^{\prime }(\mu )>0.$ Hence for
fixed $c\in \lbrack 0,2),$ the maximum of $G(\mu )$ occurs at $\mu =1.$

Also for $c=2$ we obtain%
\begin{equation}
F(\lambda ,\mu )=\frac{4}{3}\left( 1-\beta \right) ^{2}(4\beta ^{2}-8\beta
+5).  \label{eq171}
\end{equation}

Taking into account the value (\ref{eq171}), and the cases \textit{i} and 
\textit{ii}, for $0\leq \mu \leq 1$ and any fixed $c$ with $0\leq c\leq 2$,%
\begin{equation*}
\max G(\mu )=G(1)=T_{1}+T_{2}+T_{3}+T_{4}.
\end{equation*}%
For $\lambda =1$ and $0\leq \mu \leq 1$ $\left( \text{similarly }\mu =1\text{
and }0\leq \lambda \leq 1\right) ,$ we obtain%
\begin{equation*}
F(1,\mu )=H(\mu )=\left( T_{3}+T_{4}\right) \mu ^{2}+\left(
T_{2}+2T_{4}\right) \mu +T_{1}+T_{2}+T_{3}+T_{4}.
\end{equation*}

Similarly to the above cases of $T_{3}+T_{4},$ we get that%
\begin{equation*}
\max H(\mu )=H(1)=T_{1}+2T_{2}+2T_{3}+4T_{4}.
\end{equation*}

Since $G(1)\leq H(1)$ for $c\in \lbrack 0,2],$ $\max F(\lambda ,\mu )=F(1,1)$
on the boundary of the square $\mathbb{S}$. Thus the maximum of $F$ occurs
at $\lambda =1$ and $\mu =1$ in the closed square $\mathbb{S}$.

Let $K:\left[ 0,2\right] \rightarrow 
\mathbb{R}
$ 
\begin{equation}
K(c)=\max F(\lambda ,\mu )=F(1,1)=T_{1}+2T_{2}+2T_{3}+4T_{4}.  \label{eq18}
\end{equation}

Substituting the values of $T_{1},T_{2},T_{3}$ and $T_{4}$ in the function $%
K $ defined by (\ref{eq18}), yield%
\begin{equation*}
K(c)=\frac{\left( 1-\beta \right) ^{2}}{48}\left[ \left( 16\beta
^{2}-26\beta +5\right) c^{4}+24(2-\beta )c^{2}+48\right] .
\end{equation*}

Assume that $K(c)$ has a maximum value in an interior of $c\in \lbrack 0,2]$%
, by elementary calculation we find 
\begin{equation}
K^{\prime }(c)=\frac{\left( 1-\beta \right) ^{2}}{12}\left[ \left( 16\beta
^{2}-26\beta +5\right) c^{3}+12(2-\beta )c\right] .  \label{eq21}
\end{equation}

As a result of some calculations we can do the following examine:

\textbf{Case 1: }Let\textbf{\ }$16\beta ^{2}-26\beta +5\geq 0,$ that is$,$ $%
\beta \in \left[ 0,\frac{13-\sqrt{89}}{16}\right] .$ Therefore $K^{\prime
}(c)>0$ for $c\in (0,2).$ Since $K$ is an increasing function in the
interval $(0,2)$, maximum point of $K$ must be on the boundary of $c\in
\lbrack 0,2],$ that is, $c=2$. Thus, we have%
\begin{equation*}
\underset{0\leq c\leq 2}{\max }K(c)=K(2)=\frac{4}{3}\left( 1-\beta \right)
^{2}\left( 4\beta ^{2}-8\beta +5\right) .
\end{equation*}

\textbf{Case 2: }Let\textbf{\ }$16\beta ^{2}-26\beta +5<0,$ that is$,$ $%
\beta \in \left( \frac{13-\sqrt{89}}{16},1\right) .$ Then $K^{\prime }(c)=0$
implies the real critical point $c_{0_{1}}=0$ or $c_{0_{2}}=\sqrt{\frac{%
-12(2-\beta )}{16\beta ^{2}-26\beta +5}}.$ When $\beta \in \left( \frac{13-%
\sqrt{89}}{16},\frac{29-\sqrt{137}}{32}\right] ,$ we observe that $%
c_{0_{2}}\geq 2$, that is, $c_{0_{2}}$ is out of the interval $(0,2)$.
Therefore the maximum value of $K(c)$ occurs at $c_{0_{1}}=0$ or $%
c=c_{0_{2}} $ which contradicts our assumption of having the maximum value
at the interior point of $c\in \lbrack 0,2]$. Since $K$ is an increasing
function in the interval $(0,2)$, maximum point of $K$ must be on the
boundary of $c\in \lbrack 0,2],$ that is, $c=2$. Thus, we have%
\begin{equation*}
\underset{0\leq c\leq 2}{\max }K(c)=K(2)=\frac{4}{3}\left( 1-\beta \right)
^{2}\left( 4\beta ^{2}-8\beta +5\right) .
\end{equation*}

When $\beta \in \left( \frac{29-\sqrt{137}}{32},1\right) $ we observe that $%
c_{0_{2}}<2$, that is, $c_{0_{2}}$ is interior of the interval $[0,2]$.
Since $K^{\prime \prime }(c_{0_{2}})<0,$ the maximum value of $K(c)$ occurs
at $c=c_{0_{2}}.$ Thus, we have 
\begin{equation*}
\underset{0\leq c\leq 2}{\max }K(c)=K(c_{0_{2}})=K\left( \sqrt{\frac{%
-12(2-\beta )}{16\beta ^{2}-26\beta +5}}\right) =\left( 1-\beta \right)
^{2}\left( \frac{13\beta ^{2}-14\beta -7}{16\beta ^{2}-26\beta +5}\right) .
\end{equation*}

This completes the proof of the Theorem \ref{t1}.
\end{proof}

For $\beta =0$, Theorem \ref{t1} readily yields the following coefficient
estimates for bi-starlike functions.

\begin{corollary}
\label{c1} Let $f(z)$ given by (\ref{eq1}) be in the class $\mathcal{S}%
_{\sigma }^{\ast }.$ Then%
\begin{equation*}
\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert \leq \frac{20}{3}.
\end{equation*}
\end{corollary}

Our second main result for the class $\mathcal{K}_{\sigma }(\beta )$ is
following:

\begin{theorem}
\label{t2} Let $f(z)$ given by (\ref{eq1}) be in the class $\mathcal{K}%
_{\sigma }(\beta ),$ $0\leq \beta <1.$ Then%
\begin{equation}
\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert \leq \frac{\left( 1-\beta \right)
^{2}}{24}\left( \frac{5\beta ^{2}+8\beta -32}{3\beta ^{2}-3\beta -4}\right)
\label{eq22}
\end{equation}
\end{theorem}

\begin{proof}
Let $f\in \mathcal{K}_{\sigma }(\beta )\;$and $g=f^{-1}.$\ Then%
\begin{equation}
1+\frac{zf^{\prime \prime }(z)}{f^{\prime }(z)}=\beta +(1-\beta )p(z)\;\text{%
and }1+\frac{wg^{\prime \prime }(w)}{g^{\prime }(w)}=\beta +(1-\beta )q(w)
\label{eq23}
\end{equation}%
where $p(z)=1+c_{1}z+c_{2}z^{2}+...$ and $q(w)=1+d_{1}w+d_{2}w^{2}+...$ in $%
\mathcal{P}$.

Now, equating the coefficients in (\ref{eq23}), we have%
\begin{eqnarray}
2a_{2} &=&(1-\beta )c_{1},  \label{eq25} \\
6a_{3}-4a_{2}^{2} &=&(1-\beta )c_{2},  \label{eq26} \\
12a_{4}-18a_{3}a_{2}+8a_{2}^{3} &=&(1-\beta )c_{3}  \label{eq27}
\end{eqnarray}%
and%
\begin{eqnarray}
-2a_{2} &=&(1-\beta )d_{1},  \label{eq28} \\
8a_{2}^{2}-6a_{3} &=&(1-\beta )d_{2},  \label{eq29} \\
-32a_{2}^{3}+42a_{3}a_{2}-12a_{4} &=&(1-\beta )d_{3}.  \label{eq30}
\end{eqnarray}%
From (\ref{eq25}) and (\ref{eq28}), we arrive at 
\begin{equation}
c_{1}=-d_{1}  \label{eq31}
\end{equation}%
and%
\begin{equation}
a_{2}=\frac{1}{2}(1-\beta )c_{1}.  \label{eq33}
\end{equation}%
Now, from (\ref{eq26}), (\ref{eq29}) and (\ref{eq33}), we get that%
\begin{equation}
a_{3}=\frac{1}{4}\left( 1-\beta \right) ^{2}c_{1}^{2}+\frac{1}{12}\left(
1-\beta \right) \left( c_{2}-d_{2}\right) .  \label{eq34}
\end{equation}%
Also, from (\ref{eq27}) and (\ref{eq30}), we find that 
\begin{equation}
a_{4}=\frac{5}{48}\left( 1-\beta \right) ^{3}c_{1}^{3}+\frac{5}{48}\left(
1-\beta \right) ^{2}c_{1}\left( c_{2}-d_{2}\right) +\frac{1}{24}(1-\beta
)\left( c_{3}-d_{3}\right) .  \label{eq35}
\end{equation}%
Thus, we can easily establish that%
\begin{eqnarray}
&&\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert =\left\vert -\frac{1}{96}\left(
1-\beta \right) ^{4}c_{1}^{4}+\frac{1}{96}\left( 1-\beta \right)
^{3}c_{1}^{2}\left( c_{2}-d_{2}\right) \right.  \notag \\
&&\text{ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }\left. +\frac{1}{48}\left( 1-\beta
\right) ^{2}c_{1}\left( c_{3}-d_{3}\right) -\frac{1}{144}\left( 1-\beta
\right) ^{2}\left( c_{2}-d_{2}\right) ^{2}\right\vert .  \label{eq36}
\end{eqnarray}

Using (\ref{eq16}) and (\ref{eq17}) in (\ref{eq36}), we have%
\begin{eqnarray*}
&&\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert =\left\vert -\frac{1}{96}\left(
1-\beta \right) ^{4}c_{1}^{4}+\frac{1}{192}\left( 1-\beta \right)
^{3}c_{1}^{2}(4-c_{1}^{2})(x-y)\right. \\
&&+\frac{1}{48}\left( 1-\beta \right) ^{2}c_{1}\left[ \frac{c_{1}^{3}}{2}+%
\frac{(4-c_{1}^{2})c_{1}}{2}(x+y)-\frac{(4-c_{1}^{2})c_{1}}{4}(x^{2}+y^{2})+%
\frac{(4-c_{1}^{2})}{2}\left( (1-\left\vert x\right\vert
^{2})z-(1-\left\vert y\right\vert ^{2})w\right) \right] \\
&&\left. -\frac{1}{288}\left( 1-\beta \right)
^{2}(4-c_{1}^{2})^{2}(x-y)^{2}\right\vert \\
&\leq &\frac{1}{96}\left( 1-\beta \right) ^{4}c_{1}^{4}+\frac{1}{96}\left(
1-\beta \right) ^{2}c_{1}^{4}+\frac{1}{48}\left( 1-\beta \right)
^{2}c_{1}(4-c_{1}^{2}) \\
&&+\left[ \frac{1}{192}\left( 1-\beta \right) ^{3}c_{1}^{2}(4-c_{1}^{2})+%
\frac{1}{96}\left( 1-\beta \right) ^{2}c_{1}^{2}(4-c_{1}^{2})\right]
(\left\vert x\right\vert +\left\vert y\right\vert ) \\
&&+\left[ \frac{1}{192}\left( 1-\beta \right) ^{2}c_{1}^{2}(4-c_{1}^{2})-%
\frac{1}{96}\left( 1-\beta \right) ^{2}c_{1}(4-c_{1}^{2})\right] (\left\vert
x\right\vert ^{2}+\left\vert y\right\vert ^{2})+\frac{1}{576}\left( 1-\beta
\right) ^{2}(4-c_{1}^{2})^{2}(\left\vert x\right\vert +\left\vert
y\right\vert )^{2}.
\end{eqnarray*}%
Since $p\in $ $\mathcal{P},$ so $\left\vert c_{1}\right\vert \leq 2.$ Taking 
$c_{1}=c,$ we may assume without restriction that $c\in \lbrack 0,2].$ Thus,
for $\lambda =\left\vert x\right\vert \leq 1$ and $\mu =\left\vert
y\right\vert \leq 1$ we obtain%
\begin{equation}
\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert \leq M_{1}+M_{2}(\lambda +\mu
)+M_{3}(\lambda ^{2}+\mu ^{2})+M_{4}(\lambda +\mu )^{2}=\Psi (\lambda ,\mu )
\notag
\end{equation}%
where%
\begin{eqnarray*}
M_{1} &=&M_{1}(c)=\frac{\left( 1-\beta \right) ^{2}}{96}\left[ \left(
1+\left( 1-\beta \right) ^{2}\right) c^{4}-2c^{3}+8c\right] \geq 0,\text{ }
\\
M_{2} &=&M_{2}(c)=\frac{1}{192}\left( 1-\beta \right)
^{2}c^{2}(4-c^{2})(3-\beta )\geq 0, \\
M_{3} &=&M_{3}(c)=\frac{1}{192}\left( 1-\beta \right)
^{2}c(4-c^{2})(c-2)\leq 0,\text{ } \\
M_{4} &=&M_{4}(c)=\frac{1}{576}\left( 1-\beta \right)
^{2}(4-c_{1}^{2})^{2}\geq 0.
\end{eqnarray*}

Therefore we need to maximize $\Psi (\lambda ,\mu )$ in the closed square $%
\mathbb{S}=\left\{ (\lambda ,\mu ):0\leq \lambda \leq 1,0\leq \mu \leq
1\right\} .$ To show that the maximum of $\Psi $\ we can follow the maximum
of $F$ in the Theorem \ref{t1}. Thus the maximum of $\Psi $ occurs at $%
\lambda =1$ and $\mu =1$ in the closed square $\mathbb{S}$. Let $\Phi :\left[
0,2\right] \rightarrow 
\mathbb{R}
$ defined by 
\begin{equation}
\Phi (c)=\max \Psi (\lambda ,\mu )=\Psi (1,1)=M_{1}+2M_{2}+2M_{3}+4M_{4}.
\label{eq37}
\end{equation}

Substituting the values of $M_{1},M_{2},M_{3}$ and $M_{4}$ in the function $%
\Phi $ given by (\ref{eq37}), yield%
\begin{equation*}
\Phi (c)=\frac{\left( 1-\beta \right) ^{2}}{288}\left[ \left( 3\beta
^{2}-3\beta -4\right) c^{4}+4(8-3\beta )c^{2}+32\right] .
\end{equation*}

Assume that $\Phi (c)$ has a maximum value in an interior of $c\in \lbrack
0,2]$, by elementary calculation we find 
\begin{equation*}
\Phi ^{\prime }(c)=\frac{\left( 1-\beta \right) ^{2}}{72}\left[ \left(
3\beta ^{2}-3\beta -4\right) c^{3}+2(8-3\beta )c\right] .
\end{equation*}

Setting $\Phi ^{\prime }(c)=0,$ since $0<c<2,$ and $3\beta ^{2}-3\beta -4<0$
and $8-3\beta >0$ for every $\beta \in \left[ 0,1\right) $ we have the real
critical poin $c_{0_{3}}=\sqrt{\frac{2(3\beta -8)}{3\beta ^{2}-3\beta -4}}.$
Since $c_{0_{3}}\leq 2$ for every $\beta \in \left[ 0,1\right) $ and so $%
\Phi ^{\prime \prime }(c_{0_{3}})<0$, the maximum value of $\Phi (c)$
corresponds to $c=c_{0_{3}},$ that is,%
\begin{equation*}
\underset{0<c<2}{\max }\Phi (c)=\Phi (c_{0_{3}})=\Phi \left( \sqrt{\frac{%
2(3\beta -8)}{3\beta ^{2}-3\beta -4}}\right) =\frac{\left( 1-\beta \right)
^{2}}{24}\left( \frac{5\beta ^{2}+8\beta -32}{3\beta ^{2}-3\beta -4}\right) .
\end{equation*}%
On the other hand, 
\begin{equation*}
\Phi \left( 0\right) =\frac{\left( 1-\beta \right) ^{2}}{9}\text{ and }\Phi
\left( 2\right) =\frac{\left( 1-\beta \right) ^{2}}{6}\left( \beta
^{2}-2\beta +2\right) .
\end{equation*}%
Consequently, since $\Phi \left( 0\right) <\Phi \left( 2\right) \leq \Phi
(c_{0_{3}})$ we obtain$\underset{0\leq c<\leq 2}{\max }\Phi (c)=\Phi
(c_{0_{3}})$.

This completes the proof of the Theorem \ref{t2}.
\end{proof}

For $\beta =0$, Theorem \ref{t2} readily yields the following coefficient
estimates for bi-convex functions.

\begin{corollary}
\label{c2}Let $f(z)$ given by (\ref{eq1}) be in the class $\mathcal{K}%
_{\sigma }.$ Then%
\begin{equation*}
\left\vert a_{2}a_{4}-a_{3}^{2}\right\vert \leq \frac{1}{3}.
\end{equation*}
\end{corollary}

\end{document}